OpenStudy (anonymous):

f(x) = -3x - 3 g(x) = -2x^3 - 3 Find f(3) and g(-2) Thanks for any help I get :D

OpenStudy (anonymous):

f(3)=-3(3)-3=-12 g(-2)=-2(-2^3)-3=13

OpenStudy (anonymous):

can you show me how you did that? I want to learn :)